# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

### v1

@@ 5.2.2. Input Modality
Patients and attendants were enabled with three ways to interact with CataractBot: typing messages in their preferred language, recording audio, or tapping one of the three related questions suggested by the bot. However, experts could only interact with CataractBot using text. This design choice facilitated a “hygiene check” (D4) for thoroughness and precision in expert’s suggested edits, aligning with previous findings (Li et al., 2023). Out of the total 343 messages sent by patients and attendants, 206 (60.1%) were text, 43 (12.5%) audio and 94 (27.4%) related questions. The 19 participating patients sent 89 messages (43 text, 26 audio, and 20 taps), while the 30 attendants sent 245 messages (163 text, 17 audio, and 74 taps). One-way ANOVAs on message types (after Greenhouse-Geisser correction for sphericity violations) found statistical significance for attendants (F(2, 87)=13.74, p¡.001), but not for patients (F(2, 54)=0.77, p=.46). Post-hoc pairwise comparisons, corrected with Holm’s sequential Bonferroni procedure, indicated statistically significant differences between ‘text’ and ‘audio’ (t(29)=4.8, p¡.01), ‘text’ and ‘taps’ (t(29)=3.0, p¡.01), and not for ‘audio’ and ‘taps’ (t(29)=1.9, p=.11). The minimal usage of audio by attendants could be due to demographic differences, with the majority being under 45 years and well-educated. Audio messages were predominantly used by older semi-literate patients. For instance, “I can’t read (or) write Hindi that well, so asking and listening to audio, I did that.” (P12). Moreover, post-surgery limitations on screen use made speech input and audio output the preferred modality. 34 participants did not use the audio modality at all, while most participants (31) selected one of the related questions. As P7 stated: “I used the audio message once. After that there was no need for me… I just selected from the options. Click here (shows the phone) and automatically question will come, and click one question and get the answer.”. While patients and attendants found the related questions feature convenient and frequently used it, there were instances of dissatisfaction. On eight occasions, the answers to suggested questions were not available in the knowledge base, resulting in IDK responses. This issue arose because the CataractBot system use an LLM call to suggest three related questions based on the users’ last question, independent of the KB. This approach aimed to uncover gaps in the KB by not restrict questions to the curated KB, and introduce a personalized touch to the interaction by suggesting questions relevant to the ongoing conversation.
@@ 5.2.3. Multilinguality
Experts interacted with CataractBot only in English. Among our patients and attendants, the majority (29 out of 49) selected English, while 7 opted for Hindi, 6 Kannada, 5 Tamil, and 2 Telugu. None of the participants changed their initially selected language. The inclusion of four Indic languages, along with speech input functionality, ensured effective use of the bot by users with minimal literacy and tech proficiency. Given the challenges of typing in Indic languages on smartphones (Joshi et al., 2011), our non-English participants predominantly relied on speech input and selecting from related questions. Patients who chose English typed 35 questions, asked 0 questions in speech, and tapped 6 related questions. In contrast, patients opting for Indic languages typed 8 questions, sent 26 audio questions and selected 14 related questions. A similar trend was observed among attendants.
The audio queries in Indic languages were transcribed and translated to English using Azure AI services. However, due to variability in dialect across India, existing transcription and translation technologies have been found to be erroneous. It is a known issue and prior work (Jain et al., 2018) often relied on human wizards to correct errors caused by language technologies. For example, a query by P29 (Kannada) was transcribed and translated to “Krishna immediately said, how long is this kampu?”. D3, upon requesting and listening to the audio query, realized that the patient was asking about a blood clot (as ‘Kempu’ in Kannada means ‘redness’) and was able to respond appropriately. Experts faced challenged in “deciphering” such messages since they only received the English text, which was sometimes nonsensical. They suggested that they should be provided with the original audio recordings of patient queries. D4 further recommended that including audio recordings would help extract non-verbal cues (such as anxiety or the presence of other family members), enabling the expert to respond more appropriately.
It is common for people to type Indic languages in Latin script (instead of using, e.g., Devanagari script for Hindi), a feature not supported by CataractBot. We found three participants who sent a text message each by typing an Indic language query using Latin script. For one of them, the bot responded with an accurate response because the message contained a code-mix of Hindi and English:
@@
For the other two, the bot responded with “You seem to be asking the question in a language different from the selected language…” and the IDK response. Additionally, even in the selected language, participants with low literacy and language proficiency struggled to formulate questions (similar to findings in prior research (Jain et al., 2018)). For example, “Tomorrow cataract operation When should come No phone call received yet?” (P18). Such questions were common with 65 questions, of which 26 were asked using audio in an Indic language. Further analysis revealed that the speech-to-text engine was mainly responsible for these errors (19 out of 26) due to the different dialects used by the users, while the remaining 7 were translation errors.
Despite these challenges, the robustness of the GPT-4 model enabled the chatbot to answer most (32) of such ill-formulated questions appropriately. For instance, in response to the question asked by P18, the bot replied, “The surgical list is usually prepared the previous evening. If you haven’t received a call from the inpatient department, please call and confirm your timing.”
@@ 5.2.4. KB Updation
Out of the total 91 responses edited by experts, 72 (79.1%) received approval from the KB Update expert (D4) to be added to the knowledge base. In cases where the expert edits were accepted, the final response generated by the bot (by merging the initial response and expert edit suggestion) underwent additional editing by the KB Update expert in 49.3% of cases. These additional changes by D4 extended beyond information-related aspects, also addressing aspects of “tone and structure” (such as preferring active over passive voice). The changes made by the KB Update expert were minimal, with an average relative Hamming distance of 28.2%. This hints that the bot effectively incorporated the expert edit suggestions into the final bot response.
All our experts noticed an improvement in the completeness and accuracy of CataractBot’s generated answers over time, resulting in a reduction of their workload: “With time, I think the answers did get better… I am mostly just validating and saying yes.” (D3). However, there was an instance where the KB Update expert approved incorrect information (i.e., the hospital’s phone number with an extra digit) to be added to the KB. This led CataractBot provide incorrect information to three users. This error was manually corrected in the KB by a researcher on the team. Despite efforts to expand the KB, we identified limited knowledge in three medical areas—types of lenses and their comparisons, lab tests and prerequisites, and details related to prescribed medicines—and one logistical area—insurance tie-ups. Addressing it requires addition of new documents to the custom KB. Hence, D4 suggested the need for a simple way to upload documents to further expand the KB.
@@ 5.3.1. Personalization vs Privacy
In the context of evolving LLM technology, there is an ongoing discussion about privacy concerns, as LLMs have been shown to memorize training data (including personally identifiable information (PII) like emails and phone numbers) and leak it during inference (Borkar, 2023). To address these concerns, CataractBot is intentionally designed to be non-specific to individual patients, thus preserving patient privacy. A few participants (3) appreciated that their medical records were private.
@@
“Using personal health information is… violation of the patient’s personal rights. See, if that (health data) is added… external agencies can extract that information and they might use it for some advertising purposes… Personal information or any diagnosis related to any patient should not be shared.” – A17.
@@
Three educated participants (holding Master’s degrees) considered personalization as “an unnecessary add-on”. They raised concerns about potential complications associated with sharing PII data with LLMs, both in terms of legislation and “privacy threads”. A1, e.g., remarked “I would definitely not want my mom’s data to be out there, all over the internet.”. Moreover, the lack of personalization enabled users to use the chatbot for others undergoing cataract surgery: “Two three people I know are going to have their cataract surgery soon… I’ll use it for them.” (A11).
Since CataractBot neither had access to patient medical records nor was integrated with the hospital’s patient management system, it was not privy to patient-specific medical and logistical details, which resulted in IDK responses for 20 queries related to surgery schedules, updates during the procedure, and medication dosage pre-/post-surgery. During interviews, five participants expressed frustration with the IDK responses, and two of them discontinued using the chatbot after encountering a few such responses.
Participants expected personalized answers to their queries and were disappointed with generic responses. There existed an expectation mismatch, fueled by the recent hype and promises surrounding LLMs. Participants anticipated more advanced capabilities from such a system. In contrast, a participant working in the field of AI had a more balanced perspective: “See, I understand, it is just a machine… It cannot surpass or bypass or completely replace humans.” (A17). Another participant, A21, highlighted that providing more information in the chat messages could improve the bot’s ability to deliver a more personalized response. However, they acknowledged the challenge in “setting context every time”.
@@
Eight participants asked for personalized responses from the chatbot, with little concern for privacy. They felt that a personalized chatbot could save time and energy, eliminating the need to seek information from other sources or visit the hospital. This aligns with prior findings indicating minimal privacy concerns related to digital data in the global south, specifically in India (Okolo et al., 2021). Interestingly, three participants did not even consider their medical data to be “private”, as they stated, “It’s okay, it’s like nothing personal, to be hidden or something. All have common health issues.” – A14. Moreover, two participants, due to the low literacy and tech-literacy, questioned why others would be interested in their medical data, as “it is not useful for anyone else” (P18).
Educated participants were willing to share their personal medical data if the benefits outweighed the associated costs.
“If you’re traveling to Dubai and you say “I won’t show you my passport”… they will say “Thank you so much, you can’t board the plane.”… [Similarly] if you don’t want to share YOUR information [to the bot], the bot should remark “If you ask any personal questions the bot will not have access so you may have to call or visit the doctor or hospital.”… somewhere we have to compromise.” – A21.
@@
Additionally, two participants mentioned that personalization, along with the sharing of patient medical data, should be an “opt-in” feature. Individuals should conduct a cost-benefit analysis and make an informed choice based on their preferences. They believed that a personalized CataractBot would offer more relevant responses and minimize IDK responses.
@@
With each query, CataractBot provides experts with patient details such as name, age, gender, and surgery date (Figure 1A). D3 found this feature “(useful) while giving post-operative information, because (we) won’t remember patient metadata”. Four doctors and patient coordinators were in favor of integrating the chatbot with patient data, including schedules and consultation history, to offer personalized responses. Such integration would not only reduce their workload by addressing medication and logistic-related questions accurately, but it will also enhance user trust. D3 suggested collecting at least basic patient medical history to facilitate customized responses from the bot and doctors: “We don’t even know whether they’re diabetic, they’re hypertensive. So, when they ask “Should I eat and come?”… I don’t know what to tell.” (D3). Although the onboarding form attempts to capture such information in the ‘Extra Details’ field, none of the participants provide detailed information due to time constraints. C2 suggested integrating the bot with the hospitals’ EHR (Electronic Health Record) system. However, two doctors (D1 and D2) voiced concerns about CataractBot accessing patient data, citing privacy issues and skepticism about the LLMs capability to understand complex patient records accurately. “There are a lot of variables involved, including the surgeon, the particular patient, the particular case, the tests… Its too much information to process.” (D1). In addition, the KB Update expert cautioned that highly personalized responses would result in minimal updates to the KB, potentially increasing CataractBot-related workload for all experts and negatively impacting chatbot adoption among experts.
5.3.2. Bot-driven vs User-driven
CataractBot operates as a user-driven platform from the perspective of patients and attendants, while it functions as a bot-driven system for experts. Information seekers initiate conversations by asking questions and receive responses. In contrast, the bot initiates and drives verification-related conversations with experts. Patients and attendants appreciated that CataractBot “does not disturb” them, providing information solely based on the questions asked. However, a few participants suggested the bot should proactively offer relevant information about the surgery. Specifically, four attendants wished regular updates from the chatbot during the surgery, particularly while waiting anxiously outside for the procedure to conclude. Similarly, expert D4 suggested that the chatbot should notify patients: “Your discharge summary is ready, you can come and pick it up.”. A9 recommended reminders from the chatbot to ensure patients “take eye drops timely” post-surgery.
@@ 5.3.3. Reliability and Trust
Six patients/attendants explicitly mentioned trusting the bot because of their trust in the hospital, where they had chosen to undergo surgery. As CataractBot was deployed through Sankara Eye Hospital’s verified Facebook Business account, its default WhatsApp username included the hospital’s name and displayed a green tick indicating ‘verified’ status (Figures 1, 2). This contributed to participants’ trust in the bot. Moreover, the integration of the bot with the hospital was reinforced by the patient coordinator advising them to register and use the bot.
While the hospital’s reputation instilled trust in the bot, participants gained confidence in the bot’s responses due to the human experts verifying answers, denoted by the green tick icon on verified responses (Figure 1F). As A12 stated: “Basically the trust only came when we saw the green tick mark… Before that, I was also in the question mark zone… So I will say it’s a machine… But then somebody is checking it, confirming it… It’s actually then trustworthy.”. This ensured that the bot’s response were more reliable and “not fake” (P18), in contrast to a few participants who complained about receiving false information from Google searches.
Although the expert-in-the-loop feature increased trust in final responses, it had a negative impact on perceived bot intelligence. Four participants noted that every time an expert corrected CataractBot’s initial answer, trust in the bot’s unverified answers (Figure 2E) reduced. For instance, “After some point when doctors said it is invalid, I was like, okay, should I even trust the bot?” (P19). In total, 29.4% of the bot’s generated answers for medical and logistical queries were marked incorrect by the expert. This could be due to the bot’s limited custom KB or the prompting strategy favoring caution, as per doctors recommendation. A9 pointed: “It said that severe cough is dangerous on the surgery day as it increases eye pressure. But the doctor said it is completely fine… [The bot] unnecessarily alarmed me.”.
@@ 5.3.4. Accountability
CataractBot provides a persistent written record of expert-patient communication, carrying several implications. Patients mentioned utilizing the information provided by CataractBot as a reference and sharing it with others undergoing cataract surgery. Moreover, this information can be valuable when discussing their surgery with experts outside of the Sankara Eye Hospital ecosystem: “[Patients] can use the same information to be shared with their doctor at their place, “Okay, this is what was told to us.” So that [communication] becomes a little easier.” (D1). However, it also increased the sense of accountability among experts, making them “somewhat nervous”. They felt the need to be “very lawyer-like in our conversation because we don’t want it to come back and bite us tomorrow.” (D4). Furthermore, coordinators identified the system’s potential to identify logistical bottlenecks in the hospital. For example, if multiple patients are inquiring about their surgery time, it may indicate a lapse in timely communication, leading them to reach out to the bot.
@@ 6. Discussion
In this paper, we propose CataractBot, an LLM-powered WhatsApp-based expert-in-the-loop conversational agent designed to address the information needs of patients undergoing cataract surgery. Our user study revealed positive evidence of its usability among different stakeholders, including patients, attendants, doctors, and coordinators. Patients and attendants appreciated CataractBot for its ability to answer questions anytime and provide verified responses, contributing to the establishment of trust. Additionally, it streamlined communication, saving time by reducing the need for hospital calls/visits. Notably, individuals with low literacy and low tech literacy found it easy to use because of the robust multilingual and multimodal support. Experts commended CataractBot for acting as a facilitator, creating a layer between experts and patients. The asynchronous communication feature provided them the flexibility to verify LLM-generated responses at their convenience.
We discuss key design implications below.

### v2

@@ 6.2.2. Input Modality
Out of the total 343 messages sent by information seekers, 206 (60.1%) were text, 43 (12.5%) audio and 94 (27.4%) related questions. 19 participating patients sent 89 messages (43 text, 26 audio, and 20 taps), while the 30 attendants sent 254 messages (163 text, 17 audio, and 74 taps) (Figure 4C). One-way ANOVAs on message types found statistical significance for attendants (F(2, 87)=13.74, p¡.001), but not for patients (F(2, 54)=0.77, p=0.46). Post-hoc pairwise comparisons, corrected with Holm’s sequential Bonferroni procedure, indicated statistically significant differences between ‘text’ and ‘audio’ (t(29)=4.8, p¡.01) and ‘text’ and ‘taps’ (t(29)=3.0, p¡0.01) for attendants. The minimal usage of audio by attendants could be due to demographic differences, with the majority being under 45 years and well-educated. Audio messages were predominantly used by older semi-literate patients. For instance, “I can’t read (or) write Hindi that well, so asking and listening to audio, I did that.” (P12).
Post-surgery limitations on screen use also made speech input and audio output the preferred modality for patients, with 18 out of 26 audio questions by patients asked post-surgery. 34 participants did not use the audio modality at all, while 31 selected one of the related questions. As P7 stated: “I used the audio message once. After that there was no need for me… I just selected from the options. Click here (shows the phone) and automatically questions will come, and click a question and get the answer.”. While they found the related questions feature convenient and frequently used it, there were instances of dissatisfaction. On eight instances, the answers to suggested questions resulted in IDK responses. This issue arose because the CataractBot system uses an LLM call to suggest three related questions based on the users’ last question, independent of the KB. This approach aimed to uncover gaps in the KB by not restricting questions to the curated KB, and introduce personalization by suggesting questions relevant to the ongoing conversation.
With respect to experts, they could only interact with CataractBot using text. This design choice facilitated a “hygiene check” (D4) for thoroughness and precision in expert’s suggested edits, aligning with previous findings (Li et al., 2023).
@@ 6.2.3. Multilinguality
Experts interacted with CataractBot only in English. Among our patients and attendants, the majority (29 out of 49) selected English, while 7 opted for Hindi, 6 Kannada, 5 Tamil, and 2 Telugu. None of the participants changed their initially selected language. Given the challenges of typing in Indic languages on smartphones (Joshi et al., 2011), our non-English participants predominantly relied on speech input and selecting from related questions. In total, patients who chose English typed 35 questions, asked 0 questions in speech, and tapped 6 related questions. In contrast, patients opting for Indic languages typed 8 questions, sent 26 audio questions and tapped 14 related questions in total. A similar trend was observed among attendants. Patients found value in local language interaction and speech as input, helping them use the bot independently: “Without even asking me for help, he (patient) was able to use that bot as it supports Tamil… I felt it’s very useful. Because now, for elders… it is very hard to use apps and things.” – A13.
The audio queries in Indic languages were transcribed and translated to English using Azure AI services. However, this is not highly accurate due to the variability in dialect across India. It is a known issue and prior work (Jain et al., 2018) relied on human wizards to correct such errors. For example, a query by P29 (Kannada) was transcribed and translated to “I got kampu?”. D3, upon listening to the audio query, realized that the patient was asking about a blood clot (as ‘Kempu’ in Kannada means ‘redness’) and was able to respond appropriately. Experts faced challenges in “deciphering” such messages since they only received the English transcription. They suggested providing the original audio recordings of patient queries. D4 added that this would also help extract non-verbal cues, enabling them to respond more appropriately.
It is common for people to type Indic languages in Latin script (instead of using, e.g., Devanagari script for Hindi), a feature not supported by CataractBot due to current translation technology limitations. We found three participants did that. For one of them, the bot responded accurately as the message contained a code-mix of Hindi and English:
@@
For the other two, the bot responded with “You seem to be asking the question in a language different from the selected language…”. Additionally, even in the selected language, participants with low literacy and language proficiency struggled to formulate questions (similar to findings in prior research (Jain et al., 2018)). For example, “Tomorrow operation When should come No phone call received yet?” (P18). Such questions were common with 65 questions, of which 26 were asked using audio in an Indic language. Further analysis revealed that the transcription engine was mainly responsible for these errors (19 out of 26), while the remaining 7 were translation errors. Despite these challenges, the robustness of the GPT-4 model enabled the chatbot to answer most (45) of such ill-formulated questions appropriately.
@@ 6.2.4. KB Updation
Out of the total 91 responses edited by experts, 72 (79.1%) received approval from the KB update expert (D4) to be added to the KB. For 49.3% of these, the final response generated by the bot (by merging the initial response and expert edit suggestion) underwent additional editing by the KB update expert. These changes were minimal with an average relative Hamming distance of 28.2%, and mainly addressed aspects of “tone and structure”, such as preferring active over passive voice and making responses more generally applicable across patients. This hints that the bot effectively incorporated the expert edit suggestions into the final response.
Experts noticed improvements in the completeness and accuracy of CataractBot’s generated answers over time, thus reducting their workload: “With time, I think the answers did get better… I am mostly just saying yes.” (D3). However, there was an instance where the KB update expert approved incorrect information (i.e., the hospital’s phone number with an extra digit) to be added to the KB. CataractBot provided this to three seekers, before it was fixed. Despite expanding the KB, we identified our bot’s limited knowledge in two medical areas—lens types and their comparisons, and lab tests—and one logistical area—insurance tie-ups. Addressing these required manually adding new documents. D4 suggested the need for a simple way for experts to upload such documents to further expand the KB.
@@ 6.3.1. Personalization vs Privacy
Three participants appreciated that their medical records were not accessible to bot.
@@
“Using personal health information [by the bot] is… a violation of patient’s rights. If their data is added… external agencies can extract that information and might use it for advertising.” – A17.
@@
Three educated participants (holding Master’s degrees) considered personalization as “an unnecessary add-on”. They raised concerns about potential complications associated with sharing PII with LLMs, both in terms of legislation and “privacy threats”. A1, e.g., remarked “I would definitely not want my mom’s data to be out there, all over the internet.”. Moreover, lack of personalization enabled users to use the chatbot for others undergoing cataract surgery: “Two people I know are going to have their cataract surgery soon… I’ll use this bot for them.” (A11).
@@
On the other hand, fueled by the recent hype around AI and LLMs, eight participants expected “precise” and “personalized” answers to their queries with little concern for privacy, and were slightly disappointed with generic responses. This aligns with prior findings indicating minimal privacy concerns related to digital data in the global south, specifically in India (Okolo et al., 2021). Interestingly, three participants did not even consider their medical data to be “private”, as they stated, “It’s okay, it’s nothing personal, to be hidden or something. All have common health issues.” (A14). Moreover, two participants questioned why others would be interested in their medical data, as “it is not useful for anyone else” (P18).
@@
Two educated participants were willing to share their personal medical data if the benefits outweighed the costs. Additionally, two participants wanted personalization, along with the sharing of patient medical data, to be an “opt-in” feature where individuals conduct a cost-benefit analysis and make an informed choice based on their preferences. They believed that a personalized CataractBot would offer more relevant responses and minimize IDK responses.
“If you don’t want to share your information [with the bot], it should say “If you ask personal questions, I won’t have access, so you may have to visit the hospital.”… Somewhere we have to compromise.” – A21.
@@
Five experts were in favor of integrating the chatbot with patient data, including schedules and consultation history, to offer personalized responses. Such integration would not only reduce their workload, but also enhance user trust. D3 suggested collecting basic patient medical history to facilitate customized responses from the bot and doctors: “We don’t even know whether they’re diabetic, they’re hypertensive. So, when they ask “What should I eat and come?”… I ain’t sure what to tell.” (D3). Although the onboarding form attempts to capture such information in the ‘Extra Details’ field, none of our participants provided that information due to time constraints. However, D1 and D2 also voiced skepticism towards the LLMs capability to understand complex patient records accurately. “There are a lot of variables involved, including the surgeon, the particular patient, their case, tests… It’s too much information to process.” (D1). In addition, the KB update expert cautioned that highly personalized responses would result in minimal updates to the KB, potentially increasing CataractBot-related workload for experts and negatively impacting chatbot adoption among experts.
@@ 6.3.2. Reliability and Trust
Six information seekers explicitly mentioned trusting the bot because of their trust in the hospital, where they had chosen to undergo surgery. As CataractBot was deployed through Sankara Eye Hospital’s verified Facebook Business account, its default WhatsApp username included the hospital’s name and displayed a green tick indicating ‘verified’ status (Figures 1, 2). Moreover, the integration of the bot with the hospital was reinforced by the patient coordinator advising them to register and use the bot. All these contributed to participants’ trust in the bot.
With respect to bot’s responses, participants gained trust due to the human experts verifying answers (Figure 1F). As A12 stated: “Basically the trust only came when I saw the green tick mark… Before that, I was also in the question mark zone… So I will say it’s a machine… But then somebody is checking it, confirming it… It’s actually then trustworthy.”. The experts-in-the-loop ensured that the bot’s response were more reliable and “not fake” (P18). This is in contrast to a few participants who complained about receiving false information from Google search results.
Although the experts-in-the-loop feature increased trust in final responses, it had a negative impact on perceived bot intelligence. Four participants noted that when an expert corrected CataractBot’s initial answer, trust in the bot’s unverified answers reduced. For instance, “After some time, when the doctor said it is invalid, I was like, okay, should I even trust the bot?” (P19). In total, 29.4% of the bot’s generated answers for medical and logistical queries were marked incorrect by the expert. This could be attributed to the limited custom KB, the experts’ high standards for the tone and structure of answers, or our prompting strategy favoring caution. A9 pointed: “It said that severe cough is dangerous on the surgery day as it increases eye pressure. But the doctor said it is completely fine… [The bot] unnecessarily alarmed me.”
@@ 6.3.3. Accountability
CataractBot provides a persistent written record of expert-patient communication, which has implications. Patients mentioned utilizing the information provided by CataractBot as a reference and sharing it with others undergoing cataract surgery. Also, this information can be valuable when discussing their surgery with experts outside of the Sankara Eye Hospital ecosystem: “[Patients] can share the same information with their doctor at their native place… So that [communication with another doctor] becomes a little easier.” (D1). However, it also increased the sense of accountability among experts, making them “somewhat nervous”. They felt the need to be “very lawyer-like in our conversation because we don’t want it to come back and bite us tomorrow.” (D4).
@@ 7. Discussion
In this paper, we propose CataractBot, an LLM-powered WhatsApp-based experts-in-the-loop chatbot to address the information needs of patients undergoing cataract surgery. Our user study revealed positive evidence of its usability among the different stakeholders. Patients and attendants appreciated CataractBot for its ability to answer questions anytime with verified responses, which helped establish trust. Additionally, it streamlined communication, saving time by reducing the need for hospital calls/visits. Individuals with low literacy and low-tech literacy found it easy to use because of the multilingual and multimodal support. Experts commended CataractBot for acting as a facilitator, creating a layer between experts and patients. The asynchronous communication feature provided them the flexibility to verify responses at their convenience. We identify critical insights, open questions, and design implications that warrant further attention for such experts-in-the-loop chatbots to succeed on a larger scale. We discuss these aspects below.
7.1. Chatbots in Healthcare and Beyond
7.1.1. Experts-in-the-loop
Several studies have raised concerns about the application of LLMs in healthcare scenarios, due to their inconsistency, potential errors, obscured reasoning, and bias (Denecke et al., 2024; Au Yeung et al., 2023; Fogel and Kvedar, 2018; Bérubé et al., 2021; Iliashenko et al., 2019; Kocaballi et al., 2020). Moreover, users who are less (technologically) literate face a heightened risk of harm (Kapania et al., 2022). To address these concerns and comply with OpenAI’s usage policies (OpenAI, 2024), which state that “tailored medical/health advice cannot be provided without review by a qualified professional”, CataractBot incorporates experts-in-the-loop. This feature ensured reliability and fostered trust, as users knew their operating doctor was verifying the answers. However, instances where users received information from the bot that was subsequently marked as incorrect or incomplete by the expert, their trust in the initial automated answers was reduced. To address this issue, one design approach could be to eliminate the initial automated response to the patient and only send expert-verified answers. Although this might result in long wait times, which could discourage usage of the bot and prompt patients to seek information elsewhere, it is worth exploring. In the future, it would be valuable to compare these two approaches: a hybrid method of synchronous automated answers with asynchronous verification versus a fully asynchronous system with solely verified answers.
7.1.2. Conversational Design
Prior research indicates that synchronous telehealth systems can make patients feel rushed and unable to adequately express their concerns (Gordon et al., 2020; Bhat et al., 2021). In contrast, CataractBot introduces a crucial buffer in the conversation between patients and doctors, allowing them to use the system at their convenience. However, this asynchronicity also posed challenges for both parties. Patients were unsure exactly when they would receive a verified response, while experts received no further feedback from patients after providing a verification. To address this, first, patients and attendants should be made aware of the experts’ availability and informed about the expected delay in verification, particularly when questions are sent outside of working hours, on surgery days, or when experts are on leave. To achieve this, experts should be able to specify their availability in natural language on CataractBot. This will also prevent disturbing experts during their personal time. Second, to provide experts with an acknowledgement after an answer is verified, patients and attendants should be offered a poll: ‘Yes, my question was answered’—which would notify the expert, and ‘No, please give me a call’—which would alert the patient coordinator. This should be approached with caution, as prior work (Weisz et al., 2019; Yeh et al., 2022) suggests that users tend to bypass the automated agent in favour of directly connecting with a human agent via voice call.
7.1.3. Knowledge Base (KB)
For CataractBot, we curated a custom KB. Even within our short deployment, we noticed the KB to be inadequate. For longitudinal deployments, we recommend implementing a simple method for experts to add new documents directly to the KB. Recognizing the impracticality of manual updates (after approval by the KB update expert) in rapidly changing environments, we also propose automatic updates based on the consensus among doctors in the edited answers. Lastly, as the KB updates, a method is needed to evaluate CataractBot’s answers over time. Tan et al. (2024) found that GPT-4’s evaluations of AI-generated responses align closely with human clinician assessments, demonstrating the LLM’s potential not only for generating but also for validating answers. Incorporating an LLM as an evaluator has the potential to further improve the process of automatically expanding the KB while ensuring that the quality of the responses remains intact.
7.1.4. Beyond Cataract Surgery
Cataract surgery, with its brief duration, straightforward procedure, high success rate, and extensive documentation, provided an ideal low-stakes environment for building and testing our experts-in-the-loop chatbot. A similar chatbot could be even more valuable for information seekers undergoing complex, atypical, or long-term treatments, such as glaucoma surgery, cancer treatment, or childbirth. With our open-sourced framework for developing experts-in-the-loop bots, this should be feasible. Furthermore, the applicability of such a system need not be confined to the medical domain. Other fields, such as legal, finance, education, or agriculture, could benefit from similar chatbots. In these scenarios, end-users (e.g., defendants, taxpayers, students, or farmers) could receive synchronous responses from an LLM operating on a custom KB and subsequently obtain verified responses asynchronously from experts (e.g., lawyers, accountants, teachers/career counselors, or agricultural experts, respectively). Expert-mediated bots can even be used to upskill less-trained experts, such as patient coordinators or paralegals. For instance, a maternal and child health bot could serve as a valuable learning resource for community health workers in rural or semi-urban areas, who typically have minimal (medical) training. Such a bot can help them provide accurate information to patients and their families while also boosting their confidence and competence.